The mean radius of the earth is 6360 km. (a) What is the are of the earth’s surface in square meters?




Square centimeters? Square feet?

1
Expert's answer
2022-01-30T14:18:36-0500

"S=4\pi R^2,"

"S_m=5.08\cdot 10^{14}~m^2,"

"S_{cm}=5.08\cdot 10^{18}~cm^2,"

"S_{ft}=5.47\cdot 10^{15}~ft^2."